# Virtual Sensing for Real-Time Degradation Monitoring of Nuclear Systems: Leveraging DeepONet for Enhanced Sensing Coverage for Digital Twin-Enabling Technology

### v1

@@ 2.1 Data Generation
The hot leg in an AP1000 reactor plays a critical role in its operation, acting as the primary channel for transferring heat generated in the reactor core to the steam generators. Hence, the core could overheat without proper flow through the hot leg, leading to a potential meltdown. AP1000 reactor cooling system includes hot and cold leg pipes connecting the reactor vessel, steam generators, and reactor coolant pumps. Each loop has three pipes, including a 787.4 mm inner diameter pipe between the reactor vessel outlet and steam generator inlet with a length of 2.3 m Namgung and Giang (2016b). While system-level codes provide a useful tool for macroscopic thermal-hydraulic analysis of nuclear reactors, their limitations in resolving detailed flow features such as flow re-circulation and turbulence within the hot leg necessitate the use of more advanced methods like Computational Fluid Dynamics (CFD). CFD, with its ability to resolve complex geometries and flow physics, can provide a more accurate and detailed understanding of the fluid dynamics within the hot leg. However, due to the complexity of the full-scale geometry, a scaled-down model was used for CFD analysis. Geometric scaling was employed, ensuring a constant flow rate per unit volume between the actual case and the model. In this study, geometric scaling was implemented according to the following conditions:
@@
As shown in Figure 2, the elbow joint angle ($\theta_{c}$) was kept 120, same as the actual scenario of hot leg.
@@
2.2 DeepONet
In this study we have used an unstacked DeepONet described by Lu Lu et al. Lu et al. (2021), comprising a single branch network and a trunk network. Figure 4 shows the model architecture used in this work. DeepONet architecture provides a foundation for capturing the interplay between input parameters and spatial coordinates. The branch network processes the input function, representing the average initial velocity in our case. This input, belonging to an infinite-dimensional functional space, is characterized by n control points. Here n=1 as we are considering the average value. The trunk network, on the other hand, handles the spatial information. It takes as input a collection of N points within the domain, each defined by its (x, y, z) coordinates. In our study, these points correspond to 11340 virtual sensors (N=11340) distributed throughout the pipe section. The trunk network encodes the spatial dependencies inherent in the problem.
@@ 2.2 DeepONet
The crux of the DeepONet lies in the fusion of information from the branch and trunk networks. This integration is achieved through element-wise multiplication Jin et al. (2022); He et al. (2023) of the outputs from Branch and Trunk Networks. The resulting output, G(u)(y), is a function of the spatial coordinates y conditioned on the input function u. This formulation empowers the network to learn complex mappings between input and output spaces. To enhance the model’s predictive capabilities, we incorporated additional linear layers into the original DeepONet architecture. These layers enable scaling and shifting of the output, improving alignment with the target values. Figure 4 displays the original model architecture introduced by Lu Lu et al. Lu et al. (2021) alongside the modified architecture used in this study. The concept of utilizing linear layers was inspired by Kazuma et al. kobayashi2024improved_M. In the results section, we present a comparison between the models with and without the additional linear layers to further justify our modifications. The DeepONet model was trained to predict three primary output functions: turbulence, pressure, and velocity distributions in the central plane of the hot leg. These output functions reside in distinct functional spaces, denoted as $S_{1}$, $S_{2}$, and $S_{3}$, where $k\in S_{1}$, $p\in S_{2}$, and $v_{o}\in S_{3}$ represent turbulence, pressure, and velocity values in the central plane. And the mapping solution operators of the DeepONet can be defined as $G_{1}$, $G_{2}$ and $G_{3}$ for three different parameters.
@@ 3.1 Data Processing
The data for this study was generated using ANSYS Fluent, focusing on fluid dynamics within the hot leg elbow joint of an AP-1000 LWR nuclear reactor. The average inlet velocity range for the simulations was set between 0.63 and 0.83. The computational mesh consisted of a total of 11,340 nodes, distributed along the central plane of the hot leg to capture the coolant flow characteristics throughout the pipe section. For this study, we performed a total of 5,000 simulations. This dataset was subsequently divided into a training and testing set in a 90%-10% ratio, resulting in 4,500 scenarios for training and 500 scenarios for testing. Importantly, the 500 test data points were never touched during the training or validation process. They remained completely unseen by the model and were used exclusively at the final stage to evaluate the model’s performance on unseen data. This setup was designed to monitor coolant behavior in near-real-time, capturing detailed velocity, turbulence, and pressure variations to detect any anomalies effectively.
@@ 3.2 Training Process
The training process for this study is designed to ensure that the model fits the training data well while also generalizing effectively to unseen data. Initially, the training dataset is split into two parts: 80% for training and 20% for validation. The model was trained for 100 epochs to ensure convergence with this specific dataset, activation function, and optimizer. The number of hidden layers for the model was fixed at three, but the number of neurons in each layer was determined through hyperparameter tuning.
Hyperparameter tuning was achieved through Bayesian optimization using the Optuna framework, which efficiently searches for optimal parameters. The loss function employed was Mean Squared Error (MSE), and the model was optimized using the Adam optimizer with L2 regularization to prevent overfitting. Training was conducted over 100 epochs, with early stopping implemented to halt training if performance did not improve within 5 consecutive epochs, ensuring efficient use of computational resources. The model undergoes 5-fold cross-validation to assess its generalization capabilities across different subsets of the dataset. This process ensures that each data point is used for both training and validation, providing a comprehensive evaluation of the model’s performance. The average performance across all five folds gives a reliable estimate of how well the model will generalize to new data. The total training duration was 19 hours and 38 minutes. The computations for the training and evaluation tasks were performed on a computational node with a single NVIDIA A100 GPU within the Delta cluster hosted by the National Center for Supercomputing Applications (NCSA). The hyperparameter tuning process for the DeepONet neural networks involved an extensive search to identify the optimal configuration using Bayesian optimization with the Optuna framework. Table 1 summarizes the ranges and best values found for each parameter.
Table 1: Summary of Hyperparameter Tuning for DeepONet.
@@
After training the model on the entire dataset, the best model is then used to make predictions on the test dataset. The test dataset consists of data points that were not used during training or validation, providing an unbiased evaluation of the model’s performance. These 500 test data points remained completely unseen by the model and were used exclusively at this stage to evaluate the model’s performance on unseen data. To evaluate the model on test dataset following quantities were calculated:
@@
where $y_{i}$ represents the true values generated through finite element method, $\hat{y}_{i}$ represents the predicted values from nural network, and $N$ is the total number of data points.
@@ 3.3 Vanilla DeepONet vs DeepONet with Linear Layers
To support our modification of the original model with added linear layers for each parameter separately, we present a comparison between the DeepONet model with and without these linear layers. To match the total number of learnable parameters for both of the model, the number of hidden layers for Vaanilla DeepONet model was taken to be 11 for Branch network and 10 for Trunk network. Each hidden layer has 4096 neurons. Which makes the total number of learnable parameters to be 365431887. On the other hand, our considered model with linear layer has 3 hidden layers with 512 neurons each. With the extra three linear layers at the end, the total number of learnable parameters for this model is 361499727. With that these two models become comparable with each other. Table 2 shows the average value, standard deviation, and maximum value of the mean square error (MSE) and relative L2 error on the test dataset for each model, evaluated separately for the three parameters: pressure, velocity and turbulence kinetic energy.
@@ 3.5 Performance Evaluation
Parameters MSE MAE Relative L2 Error
@@
To gain a deeper understanding of the model’s predictive performance, particularly in identifying which parameters are predicted poorly and which parts of the data the model struggles with, we will include a detailed visual representation in Figure 7, 8 and 9. This figure will showcase the FEM simulated data, the neural network’s predictions, and the corresponding errors. To provide a comprehensive analysis, we have included five different plots for each parameter—pressure, velocity, and turbulence kinetic energy. These plots illustrate the best-case scenario, the worst-case scenario, and the transitional states between these extremes. By examining these visualizations, we can pinpoint the specific regions within the pipe where the model encounters difficulties in accurately predicting the coolant behavior. The best-case scenario plots highlight instances where the model’s predictions closely match the FEM simulated data, showcasing optimal performance. In contrast, the worst-case scenario plots reveal significant prediction errors with the highest discrepancies. The intermediate plots illustrate the gradual transition between these extremes, offering insights into how and where the model’s performance deteriorates. This detailed visual analysis will not only help us understand which parameters the model predicts poorly but also identify the specific sections of the pipe where these inaccuracies occur. By doing so, we can gain valuable insights into the model’s limitations and areas that require further refinement to improve the overall accuracy and reliability of the predictions.
@@ 3.6 Inference Time
DeepONet is well-known for its short inference time, especially when compared to traditional numerical methods such as the Finite Element Method (FEM). After the model has been fully trained, a well trained DeepONet model can generate predictions almost instantaneously for new inputs. This is because it bypasses the need for iterative solving of differential equations, which is computationally expensive in traditional methods like FEM.
FE Simulation Time [s] Inference Time [s]
FE Simulation 200 -
DeepONet - 1.2×$\times 10^{-3}$
Table 6: Comparison of inference time between DeepONet and FEM
Table 6 shows the inference time for DeepONet in comparison to FE simulation. As we can see, DeepONet is approximately 166,000 times faster than FEM, which requires 200 seconds for a single simulation. This drastic reduction in prediction time makes DeepONet particularly useful for virtual sensing. This further reinforces our choice of utilizing DeepONet for this study.
@@ 3.7 Local Variable Analysis and Discussion
Figure 9 represents the comparison of pressure between the CFD simulation and DeepOnet. In the 120 elbow bend, centrifugal force causes the fluid to bulge towards the convex wall, decreasing flow rate and creating a high-pressure region. Conversely, the concave wall experiences contraction effects and increased velocity, resulting in a low-pressure zone. This pressure difference between the convex and concave sides drives a secondary flow from the convex to the concave wall along the bend’s circumference. The investigation into pressure field predictions echoes previous observations regarding DeepONet’s performance in capturing velocity fields. While the model demonstrates general proficiency, it encounters challenges in regions characterized by abrupt pressure transitions, particularly near the elbow joint where steep gradients and discontinuities are prevalent. In these areas, the model exhibits elevated prediction errors compared to regions with smoother pressure distributions. This aligns with the inherent difficulty of capturing sharp gradients and discontinuities using a data-driven model like DeepONet. The pressure behavior in these zones is strongly influenced by complex flow physics, whereas DeepONet primarily learns patterns from training data.
@@
4 Conclusion
In this study, we explored the application of DeepONet as a virtual sensing framework in the context of the prediction component of a digital twin for real-time degradation monitoring of thermal-hydraulic parameters within the primary coolant loop of an AP-1000 nuclear reactor. The results demonstrate that DeepONet, especially when enhanced with additional linear layers, provides highly reliable predictions for degradation-inducing parameters (pressure, velocity, and turbulence kinetic energy) across various operating conditions. Our study also revealed that DeepONet maintains consistent performance for real-time degradation monitoring across varying percentages of training data and different numbers of data points used for training. This robustness underscores the model’s potential for deployment in scenarios with limited data or computational resources. Notably, DeepONet can make predictions 160,000 times faster than traditional finite element (FE) simulations, making it highly suited for real-time monitoring applications. However, the model faces challenges, particularly in accurately predicting turbulence kinetic energy, where it exhibited higher relative errors. This suggests that while DeepONet is a powerful tool, further refinement—possibly incorporating additional physical constraints or more sophisticated loss functions—could enhance its accuracy for more complex flow variables. Overall, this study demonstrates the feasibility and advantages of using DeepONet as a virtual sensor for monitoring coolant flow in nuclear reactors. By providing a scalable and cost-effective complement to physical sensors, DeepONet enhances the safety, reliability, and efficiency of nuclear reactors, paving the way for more advanced digital twin technologies in the nuclear industry. In future work, we will focus on improving the model’s accuracy in predicting complex flow behaviors and exploring its applicability to other critical components of nuclear reactors.

### v2

@@ 2.1 Data Generation
The hot leg in an AP1000 reactor plays a critical role in its operation, acting as the primary channel for transferring heat generated in the reactor core to the steam generators. Hence, the core could overheat without proper flow through the hot leg, leading to a potential meltdown. AP1000 reactor cooling system includes hot and cold leg pipes connecting the reactor vessel, steam generators, and reactor coolant pumps. Each loop has three pipes, including a 787.4 mm inner diameter pipe between the reactor vessel outlet and steam generator inlet with a length of 2.3 m Namgung and Giang (2016). While system-level codes provide a useful tool for macroscopic thermal-hydraulic analysis of nuclear reactors, their limitations in resolving detailed flow features such as flow re-circulation and turbulence within the hot leg necessitate the use of more advanced methods like Computational Fluid Dynamics (CFD). CFD, with its ability to resolve complex geometries and flow physics, can provide a more accurate and detailed understanding of the fluid dynamics within the hot leg. However, due to the complexity of the full-scale geometry, a scaled-down model was used for CFD analysis. Geometric scaling was employed, ensuring a constant flow rate per unit volume between the actual case and the model. In this study, geometric scaling was implemented according to the following conditions:
@@
As shown in Figure 2, the elbow joint angle ($\theta_{c}$) was kept $120^{\circ}$, same as the actual scenario of hot leg.
@@
2.2 DeepONet Architecture
In this study, we have used an unstacked DeepONet described by Lu Lu et al. Lu et al. (2021), consists of a single branch network and a trunk network. Figure 4 shows the model architecture used in this work. The DeepONet architecture lays the foundation for capturing the interplay between dynamic operational input parameters and spatially distributed system behaviors (spatial coordinates), a core characteristic of digital twin frameworks. The branch network processes the input function, which, in our case, represents the average initial velocity. This input, belonging to an infinite-dimensional functional space, is characterized by $n$ control points. Here, $n=$1, as we are considering the average value. The trunk network, on the other hand, handles the spatial information. It takes as input a collection of $N$ points within the domain, each defined by its $(x,y,z)$ coordinates. These points correspond to 11,340 virtual sensors ($N$=11,340) distributed throughout the pipe section. By encoding spatial dependencies and integrating dynamic operational data, the trunk network ensures that the digital twin remains synchronized with the physical system.
The core capability of DeepONet, which justifies its role as a digital twin, lies in the fusion of real-time information from the branch network and spatial information from trunk networks. This integration is achieved through element-wise multiplication Jin et al. (2022); He et al. (2023) of the outputs from Branch and Trunk Networks. The resulting output, $G(u)(y)$, is a function of the spatial coordinates $y$ conditioned on the input function $u$. This formulation enables the network to learn complex mappings between the input and output spaces, empowering real-time predictions of critical thermal-hydraulic parameters such as pressure, velocity, and turbulence kinetic energy. This aligns with the NRC’s definition Yadav et al. (2023) of a digital twin, which emphasizes real-time synchronization, predictive capabilities, and comprehensive system representation.
@@ 2.2 DeepONet Architecture
To enhance the model’s predictive capabilities, we added additional linear layers into the original DeepONet architecture. The concept of utilizing linear layers was inspired by the work of Kazuma et al. Kobayashi et al. (2024b). These linear layers consist of three independent layers, each dedicated to refining one of the predicted parameters - velocity, turbulence kinetic energy, or pressure. Each linear layer takes an input vector corresponding to the spatial domain (in our case 11,340 spatial points) and outputs a vector of the same size, ensuring that the predictions retain the spatial resolution required for monitoring. The weights of these layers are initialized using Xavier initialization, which ensures stable gradients and efficient training, while biases are initialized to zero to prevent unnecessary offsets in the early stages. This modular design allows each linear layer to specialize in its respective parameter, enabling parameter-specific transformations that improve prediction accuracy. These layers enable scaling and shifting of the output, improving alignment with the target values. Figure 4 displays the original model architecture introduced by Lu Lu et al. Lu et al. (2021) (a), alongside the modified architecture used in this study (b). After the element-wise multiplication of the outputs from the Branch and Trunk Networks, the resulting tensor is split into three separate components, with each component corresponding to a specific parameter (velocity, turbulence, or pressure). Each component is then passed through its respective linear layer, which applies a final transformation. For a batch of data, the input tensor to the linear layers has dimensions (batch-size, no-of-parameters, number-of-spatial-points). After processing, each linear layer outputs a tensor of size (batch-size, number-of-spatial-points), in our case $(512,11340)$, preserving spatial resolution by mapping input vectors directly to corresponding output vectors. In section 3.3, we present a comparison between the models with and without the additional linear layers to further justify this modifications. The DeepONet model was trained to predict three primary output functions: turbulence, pressure, and velocity distributions in the central plane of the hot leg. These output functions reside in distinct functional spaces, denoted as $S_{1}$, $S_{2}$, and $S_{3}$, where $k\in S_{1}$, $p\in S_{2}$, and $v_{o}\in S_{3}$ represent turbulence, pressure, and velocity values in the central plane. And the mapping solution operators of the DeepONet can be defined as $G_{1}$, $G_{2}$, and $G_{3}$ for three different parameters.
@@ 3.1 Data Processing
The data for this study was generated using ANSYS Fluent, focusing on fluid dynamics within the hot leg elbow joint of an AP-1000 LWR nuclear reactor. The average inlet velocity range for the simulations was set between 0.63 and 0.83. The computational mesh consisted of a total of 11,340 nodes, distributed along the central plane of the hot leg to capture the coolant flow characteristics throughout the pipe section. For this study, we performed a total of 5,000 simulations. This dataset was subsequently divided into a training and testing set in a 80%-20% ratio, resulting in 4,000 scenarios for training and 1000 scenarios for testing. Importantly, the 1000 test data points were never touched during the training or validation process. They remained completely unseen by the model and were used exclusively at the final stage to evaluate the model’s performance on unseen data. This setup was designed to monitor coolant behavior in near-real-time, capturing detailed velocity, turbulence, and pressure variations to detect any anomalies effectively.
@@ 3.2 Training Process
The training process for this study is designed to ensure that the model fits the training data well while also generalizing effectively to unseen data. Initially, the training dataset is split into two parts: 80% for training and 20% for validation. The final model, including the linear layer, was trained for 100 epochs to ensure convergence with this specific dataset, activation function, and optimizer. The number of hidden layers for the final model was fixed at three, but the number of neurons in each layer was determined through hyperparameter tuning.
Hyperparameter tuning for the final model with linear layers was achieved through Bayesian optimization using the Optuna framework, which efficiently searches for optimal parameters. The loss function employed was Mean Squared Error (MSE), and the model was optimized using the Adam optimizer with L2 regularization to prevent overfitting. Training was conducted over 100 epochs, with early stopping implemented to halt training if performance did not improve within 5 consecutive epochs, ensuring efficient use of computational resources. The model underwent 5-fold cross-validation to assess its generalization capabilities across different subsets of the dataset. This process ensures that each data point is used for both training and validation, providing a comprehensive evaluation of the model’s performance. The average performance across all five folds gives a reliable estimate of how well the model will generalize to new data. The total training duration was 19 hours and 38 minutes. The computations for the training and evaluation tasks were performed on a computational node with a single NVIDIA A100 GPU within the Delta cluster hosted by the National Center for Supercomputing Applications (NCSA). While hyperparameter tuning determined most parameters, the number of linear layers was decided through manual inspection based on observed improvements in performance. Table 1 summarizes the ranges and best values found for each parameter.
Table 1: Summary of Hyperparameter Tuning for DeepONet with linear layers.
@@
After training the model on the entire dataset, the best model is then used to make predictions on the test dataset. The test dataset consists of data points that were not used during training or validation, providing an unbiased evaluation of the model’s performance. These 1000 test data points remained completely unseen by the model and were used exclusively at this stage to evaluate the model’s performance on unseen data. To evaluate the model on test dataset following quantities were calculated:
@@
where $y_{i}$ represents the true values generated through finite volume method, $\hat{y}_{i}$ represents the predicted values from neural network, and $N$ is the total number of data points.
The values reported in the tables are calculated as the averages over the 1000 test scenarios in the test set, ensuring an unbiased evaluation of model performance. The average values for Mean Square Error (MSE), Mean Absolute Error (MAE), and Relative L2 Error are computed as follows:
Average $\text{Average Metric}=\frac{1}{M}\sum_{j=1}^{N}\text{Metric}_{j}$ (6)
Here, $\text{Metric}_{j}$ is the value of the metric (MSE, Relative L2 Error, MAE) for the $j$-th test scenario. And $M$ is total number of test scenarios (1000 in this case). The formula aggregates the metric for all test scenarios and divides it by the total number of scenarios to compute the average.
@@ 3.3 Vanilla DeepONet vs DeepONet with Linear Layers
To support our modification of the original model with added linear layers for each parameter separately, we present a comparison between the DeepONet model with and without these linear layers. This comparison is conducted purely to evaluate the impact of adding linear layers, and the Vanilla DeepONet model is used solely for this purpose.
To match the total number of learnable parameters for both of the model, the number of hidden layers for Vanilla DeepONet model was taken to be 11 for Branch network and 10 for Trunk network. Each hidden layer has 4096 neurons. Which makes the total number of learnable parameters to be 365431887. On the other hand, our considered model with linear layer has 3 hidden layers with 512 neurons each. With the extra three linear layers at the end, the total number of learnable parameters for this model is 361499727. With that these two models become comparable with each other. Table 2 shows the average value, standard deviation, and maximum value of the mean square error (MSE) and relative L2 error on the test dataset for each model, evaluated separately for the three parameters: pressure, velocity and turbulence kinetic energy. It is important to note that no hyperparameter tuning was performed for the Vanilla DeepONet model. The parameters for this model were taken to match those of the DeepONet model with linear layers, as described in Section 3.2. Since the Vanilla DeepONet model is not used for further evaluation or visualization, hyperparameter tuning was not necessary.
@@ 3.5 Performance Evaluation
Parameters Average MSE Average MAE Average Relative L2 Error
@@
A detailed visual analysis of the model’s predictive performance, including best-case, worst-case, and transitional scenarios, is presented in Section 3.7 to further investigate the model’s strengths and limitations.
@@ 3.6 Inference Time
DeepONet is well-known for its short inference time, especially when compared to traditional numerical methods such as the Finite Volume Method (FVM). After the model has been fully trained, a well trained DeepONet model can generate predictions almost instantaneously for new inputs. This is because it bypasses the need for iterative solving of differential equations, which is computationally expensive in traditional methods like FVM.
FV Simulation Time [s] Inference Time [s]
FV Simulation 200 -
DeepONet - 1.3507×$\times 10^{-1}$
Table 6: Comparison of inference time between DeepONet and FVM
Table 6 shows the inference time for DeepONet in comparison to FVM simulation. As we can see, DeepONet is approximately 1481 times faster than FVM, which requires 200 seconds for a single simulation. This drastic reduction in prediction time makes DeepONet particularly useful for virtual sensing. This further reinforces our choice of utilizing DeepONet for this study.
Note: Both the FV simulation and DeepONet inference were performed on the same machine to ensure a fair comparison. The machine specifications were Intel Core i7 CPU with 16 GB RAM.
@@ 3.7 Local Variable Analysis and Discussion
To provide a deeper understanding of the model’s predictive performance, we include a comprehensive visual analysis in Figure 7, 8 and 9. These figures showcase the ANSYS simulated data, the neural network’s predictions, and the corresponding errors. For each parameter—pressure, velocity, and turbulence kinetic energy—we include five plots: the best-case scenario, the worst-case scenario, and three intermediate cases, all selected based on Relative $L2$ Error. The best-case plots highlight scenarios where the model’s predictions closely align with FVM data, demonstrating optimal performance. Conversely, the worst-case plots reveal significant discrepancies and prediction errors. The intermediate plots show the gradual transition between these extremes, providing insights into regions where the model’s performance deteriorates. This detailed visual analysis will not only help us understand which parameters the model predicts poorly but also identify the specific sections of the pipe where these inaccuracies occur. We can gain valuable insights into the model’s limitations and highlight areas for refinement to enhance its overall accuracy and reliability.
@@
Figure 9 represents the comparison of pressure between the CFD simulation and DeepOnet. In the $120^{\circ}$ elbow bend, centrifugal force causes the fluid to bulge towards the convex wall, decreasing flow rate and creating a high-pressure region. Conversely, the concave wall experiences contraction effects and increased velocity, resulting in a low-pressure zone. This pressure difference between the convex and concave sides drives a secondary flow from the convex to the concave wall along the bend’s circumference. The investigation into pressure field predictions echoes previous observations regarding DeepONet’s performance in capturing velocity fields. While the model demonstrates general proficiency, it encounters challenges in regions characterized by abrupt pressure transitions, particularly near the elbow joint where steep gradients and discontinuities are prevalent. In these areas, the model exhibits elevated prediction errors compared to regions with smoother pressure distributions. This aligns with the inherent difficulty of capturing sharp gradients and discontinuities using a data-driven model like DeepONet. The pressure behavior in these zones is strongly influenced by complex flow physics, whereas DeepONet primarily learns patterns from training data.
@@
4 Discussion
In this study, we have explored the integration of Digital Twin technology with advanced neural operator models, specifically DeepONet, to develop a virtual sensing-enabled framework for real-time parameter prediction in nuclear reactors. This section elaborates on the significance of our findings, the role of DeepONet within a Digital Twin framework, and the implications of our results for degradation monitoring. Additionally, we address the limitations of our approach and discuss potential future directions.
4.1 DeepONet-based Digital Twin
The DeepONet developed in this study serves as a digital twin framework for plant operations, aligning with the U.S. Nuclear Regulatory Commission’s (NRC’s) definition of a digital twin through its integration of real-time inference, adaptability to operational condition changes, and synchronization with the physical system Yadav et al. (2023); Liu et al. (2024).
  • Dynamic Digital Representation and Synchronization with the Physical System: DeepONet provides a dynamic representation of the system it models, offering real-time insights into the system’s behavior. Synchronization between the physical system and its digital representation is a key requirement for realizing digital twin technology. DeepONet addresses this need by processing real-time signals through a supporting data pipeline and dynamically updating its predictions, enabling seamless interaction with the physical system Liu et al. (2024). DeepONet models can be trained using simulation data (such as ANSYS simulations used in this study) and later fine-tuned with real-world data, bridging the gap between digital and physical systems. This capability aligns with the key feature of digital twins to adapt and evolve alongside their physical counterparts. As shown in our study, the branch network processes real-time sensor data (average inlet velocity) ensuring synchronization, while the trunk network provides spatial details, functioning as a virtual sensor network. This setup allows DeepONet to augment physical sensor data by providing full-field predictions, such as flow velocity and turbulence, at critical locations like pipe bends and elbow joints, where physical sensors are impractical. The model’s predictive capabilities, as shown in Figures 7, 8, and 9, demonstrate its effectiveness in replicating system behavior. These results validate DeepONet’s role as a reliable and efficient component of the Digital Twin framework.
  • Real-Time Inference: A key feature of digital twins is real-time or near-real-time inference to provide operational insights and enhance decision-making during dynamic conditions. As shown in Table 6, DeepONet achieves unprecedented speed, making predictions 1481 times faster than traditional finite volume simulations. This allows operators to monitor critical thermal-hydraulic parameters—pressure, velocity, and turbulence kinetic energy—almost instantaneously. The real-time inference capability ensures the model remains an active, dynamic component of the digital twin framework, crucial for nuclear plant operations where quick decision-making is essential.
  • Adaptability to Operational Condition Changes: Digital twins must adapt to changing operational conditions and maintain accuracy across diverse scenarios without constant retraining. This study demonstrates that DeepONet fulfills this requirement by learning the underlying functional relationships between input parameters and spatial coordinates. Its architecture ensures robust predictions under varying conditions, without requiring retraining. As shown in Figure 6, DeepONet performs consistently across diverse scenarios. It effectively predicts spatial distributions of velocity, turbulence, and pressure across different inlet conditions, showcasing its adaptability and reliability for monitoring dynamic systems. This capability aligns with the NRC’s definition of digital twins, emphasizing adaptability and synchronization with evolving physical systems.
These findings validate DeepONet’s role in enabling a Digital Twin framework, bridging real-time data and predictive modeling.
4.2 Monitoring Operational Conditions Indicative of Degradation
Degradation, such as material/wall thinning, stress corrosion cracking, or fatigue, develops gradually due to persistent mechanical stresses and flow irregularities. Certain operational conditions, however, may act as early indicators of these processes and can be monitored to prevent long-term damage. For instance, flow-induced vibrations, caused by turbulent flow around pipe bends and fittings, are a major contributor to fatigue and erosion-corrosion in nuclear piping systems. These vibrations can lead to high-frequency stresses, overloading pipe supports or nozzles and accelerating material degradation. Similarly, transient phenomena such as water hammer or condensation-induced hammer, generate impact loads with high dynamic factors, which can weaken pipe walls or connections over time. Monitoring turbulence pattern, pressure fluctuations, and velocity distributions in such areas provides actionable insights into regions where degradation risks are elevated. Jacimovic and D’Agaro (2020)
While our study does not directly simulate material degradation or wall thinning, the ability to monitor turbulence and pressure in areas prone to stress highlights DeepONet’s usefulness in degradation-informed decision-making. Elevated turbulence near pipe bends is a known precursor to erosion-corrosion, while fluctuating pressure often signals transient phenomena such as water hammer, both of which can severely impact system integrity. So say for example, a rise in turbulence intensity near elbow joints (Figure 8) over time could signal heightened mechanical stress in those regions, suggesting targeted inspections or preventive maintenance. Similarly, deviations in pressure (as in Figure 9) or velocity distributions (shown in Figure 7) can identify developing flow-induced vibrations that may accelerate wear, particularly in locations vulnerable to acoustic resonance.
Within a Digital Twin, these data can be collected over time to form flow profiles, revealing how conditions within a pipe evolve dynamically. Such profiles are invaluable for identifying patterns indicative of emerging degradation risks, such as increasing pressure in certain region or abnormal turbulence patterns near pipe bends. Furthermore, by illuminating unmonitored regions with accurate predictions, DeepONet enhances the visibility of the entire system, enabling reliable condition-based maintenance strategies, that can reduce maintenance cost and enhance reactor safety.
4.3 Limitations and Future Directions
While the results of this study show the effectiveness of DeepONet in predicting key thermal-hydraulic parameters, certain limitations must be acknowledged. One notable challenge is the spectral bias inherent in data-driven models like DeepONet. All Neural networks tend to prioritize learning low-frequency, smoother patterns over high-frequency, more complex ones, such as those observed in turbulent regions. This spectral bias, coupled with the imbalanced nature of our dataset, evident in Figures 7, 8, and 9, where larger, smoother flow patterns dominate over the relatively smaller turbulent flow patterns in the bend region, likely contributed to the increased error observed in predicting turbulence. We assume that the model is smoothing out high-frequency patterns. For example, the model exhibited higher errors in areas with intense turbulence, such as near elbow joints and bends, where high-frequency flow patterns dominate. These regions are most important for identifying flow-induced vibrations and other precursors to degradation, highlighting a potential limitation of the current approach.
To address this, future research could explore combining neral operator with diffusion models Oommen et al. (2024). Future directions could also include designing hybrid frameworks, where specialized models can be trained for different regions of the system based on the initial CFD modeling. This targeted approach could improve the performance by allowing each model to focus on specific flow characteristics, thereby addressing challenges such as spectral bias and ensuring better accuracy in regions with high-frequency turbulent flows.
Acknowledgment
The authors were supported by the National Science Foundation (award OAC 2005572 for Delta supercomputing facility), the United States Department of Energy (DOE), and Battelle Energy Alliance LLC (Idaho National Lab).